### v3

@@ 3.4 Ne/O Ratios
As shown in the lower right panels of FigureĀ 4, the 10,000 $M_{\odot}$Ā star with 10% mass loss is the best match to GS 3073, with Ne/O == -0.71 at a dilution factor of $\sim$ 20. Only one model here produces an overabundance of Ne/O relative to GS 3073. For 50% mass loss the 9000 $M_{\odot}$Ā model is the best match, with Ne/O == -0.77 at a dilution factor of $\sim$ 20. Two of the 10 models produce an overabundance of Ne/O. Finally, for the CO core case two stars match GS 3073, with the 2000 $M_{\odot}$Ā model coming closest with Ne/O == -0.77 at a dilution factor of $\sim$ 20. Four models here produce an overabundance of Ne/O.
3.5 Stars Below 1000 $M_{\odot}$Ā or Above 10,000 $M_{\odot}$
We find that N/O ratios for GS 3073 cannot be explained by Pop III SMSs above 10,000 $M_{\odot}$Ā or below 1000 $M_{\odot}$. 900 and 950 $M_{\odot}$Ā stars evolved until the end of core O burning do not produce N/O ratios greater than -0.16, regardless of mass loss, and 15,000 $M_{\odot}$Ā and 21,000 $M_{\odot}$Ā stars evolved to the end of core C and He burning produce an N/O of at most 0.19, still well below the observed 0.46. The 21,000 $M_{\odot}$Ā star also has a much lower O/H ++12 ratio of 3 - 6, effectively ruling it out as a candidate. These stars strongly suggest the presence of an upper and lower limit in SMS mass beyond which the GS 3073 N/O ratio cannot be produced.
@@ 4 Discussion and Conclusion
Figure 5: Abundance ratios of 10 SMSs with contributions by the galactic stars for the case of 10% mass loss. If the chemically mature stellar population contributes C, N, O, and Ne mass fractions $\chi_{\rm C}=$ $2.8\times 10^{-4}$, $\chi_{\rm N}=$ $8\times 10^{-5}$, $\chi_{\rm O}=$ $1.2\times 10^{-3}$ and $\chi_{\rm Ne}=$ $9.4\times 10^{-4}$ then all 10 models can reproduce the observed near vertical transition in N/O ratio while also explaining the C/O ratios (Ji etĀ al., 2024).
@@
No single SMS would be expected to match all the abundance ratios of GS 3073 because other generations of stars also contributed to them. Their contribution decreased N/O ratios and elevated C/O ratios over time because mass loss and SNe produced much more C and O than N (Ekström et al., 2012; Murphy et al., 2021). SMSs with N/O ratios greater than those in GS 3073 and C/O ratios less than those in GS 3073 can thus still be the origin of its N/O excess and C/O fractions. We can estimate the contribution to N/O, C/O and Ne/O ratios in GS 3073 from its less massive stars by comparison to surface abundances of halo and disk stars in the MW at similar metallicities (Prantzos et al., 2018). From Figures 4 and 8 of Ji et al. (2024) we find N and O mass fractions for this population to be 8.0e-5 and 1.2e-3, respectively. Assuming [Fe/H] $\sim$ -0.6 in GS 3073 (Ji et al., 2024), we then obtain C/Fe and Ne/Fe ratios and then C and Ne abundances of 2.8e-4 and 9.4e-4, respectively. We then mix ejecta from the SMS with the ISM for a range of dilution factors.
@@
As shown in FigureĀ 5, when the SMS and stellar contributions to N/O, C/O, and Ne/O are taken together, we obtain good agreement with those of GS 3073 with a 6000 $M_{\odot}$Ā star with 10% mass loss in this scenario. Mixing of the SMS and ISM components produces a nearly vertical shift in N/O versus O/H from its floor of -1.1 to 0.46, suggesting the significant increase in N abundance without a substantial change in O abundance shown in Figure 5 of Ji etĀ al. (2024). This test also shows that the SMSs must be Pop III because no other stars would keep O/H constant. Ours are also the only known stars that can produce the Ne/O ratio in GS 3073. While GR supernovae of metal-enriched SMSs can explain the smaller N excess in GN-z11 (Nagele & Umeda, 2023), they cannot account for its Ne abundances. GR SNe also blow most of the gas from their host galaxies, greatly delaying subsequent star formation (Whalen etĀ al., 2013a, b, d; Johnson etĀ al., 2013; Chen etĀ al., 2014a), so it is not clear if GS 3073 could have reached its stellar mass, 109.5 $M_{\odot}$, by $z=$ 5.55 if a GR SN had occurred early in its history. Furthermore, such an event cannot explain the presence of a BH in GS 3073 because it would not have created one, whereas an SMS can account for both the nitrogen excess in GS 3073 and its AGN. Not only do several of our stars match Ne/O ratios for GS 3073, some produce an overabundance of Ne that would be diluted down to the observed values by other populations of stars.
@@
Gas at lower densities in GS 3073, as traced by optical lines, has a much lower N/O == -1.1 than gas at high densities, which is traced by UV lines (Ji etĀ al., 2024). Stratified N/O ratios are consistent with our SMS scenario, as cosmological simulations indicate that SMSs form in dense gas in high-redshift halos. Whether the SMS formed in a satellite halo and later merged with GS 3073 in which a mature population of stars had already formed (Regan etĀ al., 2017) or it formed in a pristine pocket of gas in GS 3073 (Liu & Bromm, 2020; Venditti etĀ al., 2023), there would have been the near vertical rise in N/O versus O/H. Because the true distribution of low-metallicity and high-metallicity gas in GS 3073 is not known, it is not yet possible to determine precisely which enrichment scenario happened, only that massive nitrogen production by SMSs must have occurred.
@@
DN was supported by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No 833925, project STAREX) and by the Virginia Initiative of Theoretical Astronomy (VITA) Fellowship. AH acknowledges support by the Australian Research Council (ARC) through grants DP240101786 and DP240103174. MAL was funded by UAEU via UPAR grant No. 12S111.

### v4

@@ 3.4 Ne/O Ratios
As shown in the lower right panels of Figure 4, the 10,000 $M_{\odot}$ star with 10% mass loss is the best match to GS 3073, with Ne/O == -0.71 at a dilution factor of $\sim$ 20. Only one model here produces an overabundance of Ne/O relative to GS 3073. For 50% mass loss the 9000 $M_{\odot}$ model is the best match, with Ne/O == -0.77 at a dilution factor of $\sim$ 20. Two of the 10 models produce an overabundance of Ne/O. Finally, for the CO core case two stars match GS 3073, with the 2000 $M_{\odot}$ model coming closest with Ne/O == -0.77 at a dilution factor of $\sim$ 20. Four models here produce an overabundance of Ne/O.
3.5 Stars Below 1000 $M_{\odot}$ or Above 10,000 $M_{\odot}$
We find that N/O ratios for GS 3073 cannot be explained by Pop III SMSs above 10,000 $M_{\odot}$ or below 1000 $M_{\odot}$. 900 and 950 $M_{\odot}$ stars evolved until the end of core O burning do not produce N/O ratios greater than -0.16, regardless of mass loss, and 15,000 $M_{\odot}$ and 21,000 $M_{\odot}$ stars evolved to the end of core C and He burning produce an N/O of at most 0.19, still well below the observed 0.46. The 21,000 $M_{\odot}$ star also has a much lower O/H ++12 ratio of 3 - 6, effectively ruling it out as a candidate. These stars strongly suggest the presence of an upper and lower limit in SMS mass beyond which the GS 3073 N/O ratio cannot be produced.
The emergence of this window is also captured by the integrated ejecta shown in Table 1. Defining
$M_{\rm N}^{\rm ej}$ $=\int_{m_{\rm cut}}^{M}X_{\rm N}(m)\,\mathrm{d}m,$
$M_{\rm O}^{\rm ej}$ $=\int_{m_{\rm cut}}^{M}X_{\rm O}(m)\,\mathrm{d}m.$
[…]
@@ 4 Discussion and Conclusion
Figure 5: Abundance ratios of 5 SMSs with contributions by the galactic stars for the case of 10% mass loss. The green, yellow and brown stars indicate corresponding ratios for GS 3073, CEERS 1019 and GN-z11, respectively. If the chemically mature stellar population contributes C, N, O, and Ne mass fractions $\chi_{\rm C}=$ $2.8\times 10^{-4}$, $\chi_{\rm N}=$ $8\times 10^{-5}$, $\chi_{\rm O}=$ $1.2\times 10^{-3}$ and $\chi_{\rm Ne}=$ $9.4\times 10^{-4}$ then all 10 models can reproduce the observed near vertical transition in N/O ratio while also explaining the C/O ratios (Ji et al., 2024).
@@
No single SMS would be expected to match all the abundance ratios of GS 3073 because other generations of stars also contributed to them. Their contribution decreased N/O ratios and elevated C/O ratios over time because mass loss and SNe produced much more C and O than N (Ekström et al., 2012; Murphy et al., 2021). SMSs with N/O ratios greater than those in GS 3073 and C/O ratios less than those in GS 3073 can thus still be the origin of its N/O excess and C/O fractions. We can estimate the contribution to N/O, C/O and Ne/O ratios in GS 3073 from its less massive stars by comparison to surface abundances of halo and disk stars in the MW at similar metallicities (Prantzos et al., 2018). From Figures 4 and 8 of Ji et al. (2024) we find N and O mass fractions for this population to be 8.0e-5 and 1.2e-3, respectively. Assuming [Fe/H] $\sim$ -0.6 in GS 3073 (Ji et al., 2024), we then obtain C/Fe and Ne/Fe ratios and then C and Ne abundances of 2.8e-4 and 9.4e-4, respectively. We then mix ejecta from the SMS with the ISM for a range of dilution factors. The present grid differs from Nandal et al. (2024b) by advancing to core–O (and in some cases core–Si) burning with TOV corrections, by adding Ne/O as a key abundance test, by confronting the tighter abundance set of GS 3073, and by including an exploratory rotating SMS subset to assess sensitivity to rotational mixing during He burning. Additionally, we also explain the black hole mass of GS3073 by invoking sub-Eddington accretion onto the seed formed by such SMSs at redshifts up to 20.
@@
As shown in Figure 5, when the SMS and stellar contributions to N/O, C/O, and Ne/O are taken together, we obtain good agreement with those of GS 3073 with a 6000 $M_{\odot}$ star with 10% mass loss in this scenario. Mixing of the SMS and ISM components produces a nearly vertical shift in N/O versus O/H from its floor of -1.1 to 0.46, suggesting the significant increase in N abundance without a substantial change in O abundance shown in Figure 5 of Ji et al. (2024). This test also shows that the SMSs must be Pop III because no other stars would keep O/H constant. Ours are also the only known stars that can produce the Ne/O ratio in GS 3073. While GR supernovae of metal-enriched SMSs can explain the smaller N excess in GN-z11 (Nagele & Umeda, 2023), they cannot account for its Ne abundances. GR SNe also blow most of the gas from their host galaxies, greatly delaying subsequent star formation (Whalen et al., 2013a, b, d; Johnson et al., 2013; Chen et al., 2014a), so it is not clear if GS 3073 could have reached its stellar mass, 109.5 $M_{\odot}$, by $z=$ 5.55 if a GR SN had occurred early in its history. Furthermore, such an event cannot explain the presence of a BH in GS 3073 because it would not have created one, whereas an SMS can account for both the nitrogen excess in GS 3073 and its AGN. Not only do several of our stars match Ne/O ratios for GS 3073, some produce an overabundance of Ne that would be diluted down to the observed values by other populations of stars.
@@
Gas at lower densities in GS 3073, as traced by optical lines, has a much lower N/O == -1.1 than gas at high densities, which is traced by UV lines (Ji et al., 2024). Stratified N/O ratios are consistent with our SMS scenario, as cosmological simulations indicate that SMSs form in dense gas in high-redshift halos. Whether the SMS formed in a satellite halo and later merged with GS 3073 in which a mature population of stars had already formed (Regan et al., 2017) or it formed in a pristine pocket of gas in GS 3073 (Liu & Bromm, 2020; Venditti et al., 2023), there would have been the near vertical rise in N/O versus O/H. Because the true distribution of low-metallicity and high-metallicity gas in GS 3073 is not known, it is not yet possible to determine precisely which enrichment scenario happened, only that massive nitrogen production by SMSs must have occurred.
@@
The authors would like to thank the referee for their invaluable comments and feedback on the manuscript. DN was supported by the Swiss National Science Fund (SNSF) Postdoctoral Fellowship, grant number: P500-2235464 and by the Virginia Initiative of Theoretical Astronomy (VITA) Fellowship. AH acknowledges support by the Australian Research Council (ARC) through grants DP240101786 and DP240103174. MAL was funded by UAEU via UPAR grant No. 12S111.